OpenStudy (anonymous):

The Mogul Runners ski club planned a trip to Park City. Of the total number of club members, 11 signed up to go. If this is 55% of the club, how many members does the ski club have?

OpenStudy (anonymous):

Total number of members signed = 11 It is said that this is 55% of the total members of the club Lets suppose total members as x now it is 55/100*x = 11 55x = 1100 x = 1100/55 = 20

OpenStudy (anonymous):

Find the measure of ∠AED for m∠BEC = 96 degrees